what is the answer?(1 point) Finding the length of a curve. Arc length for y = f(x). Let f(x) be a smooth function over the interval [a, b]. The
Answer #1

partia fox = 2 lnoo - x² >F6-da (2600 ] on 16 2 In 1 =- = 2.5 to (20) = ? ano 5F(G)- E-) - BD2+(32-2- = 3)² + ( 2 ) 2 - + L (L- 64x² + x4 +256-320² (80) 2 J x4 +256 +32x2 - dx 5 (88) (x2 +16) 2 da 80 Hit - ST (22+19) de ada = lnalec * +16. Antal 177